OpenStudy (anonymous):

Find the linear function f with f -1 (-8) = 5 and f -1 (-4) = -5.

OpenStudy (anonymous):

u meant f inverse by -1 rite?

OpenStudy (anonymous):

yes, inverse

OpenStudy (anonymous):

so for finding the inverse function of function inverse, just reflect the given function wrt y=x line. means u need to find the line joining (5,-8) and (-5,-4).

OpenStudy (anonymous):

do you find that by simply creating the line using those 2 pts?

OpenStudy (anonymous):

I keep getting 0 for the gradient/slope?

OpenStudy (anonymous):

yes , the slope is (-4-(-8))/(-5-5)=-2/5

OpenStudy (anonymous):

aaag; I was calcing Y^2 - man Im rusty - thanks a million!

OpenStudy (anonymous):

y=-2/5x-6

OpenStudy (anonymous):

done :)